arshad bought 50 shares for 60 each, and after months he bought 25 shares for 56 each, then at what price arshad must buy more 25 shares, as the average price become 58.
Answer by mananth(16946)

arshad bought 50 shares for 60 each, = 3000
25 shares for 56 each,= 1400
25 shares, x
Average =58
Total value = 58*100 = 5800
50*60+56*25+25x= 5800
3000+1400+25x=5800
25x=5800-4400
25x=1400
/25
x=56